OpenStudy (anonymous):

which equation represents a linear function? A.)y=-x^2-4 B.)-3x^2+1=y C.)y=x^2 D.)x+2=y

OpenStudy (anonymous):

a first degree polynomial. so the x has a power of 1

OpenStudy (anonymous):

constant slope

OpenStudy (anonymous):

mmmmmm still confused on this question??....

OpenStudy (anonymous):

A linear function will be in the format y=mx+b

OpenStudy (anonymous):

yes, that is correct. The highest degree of polynomial for a linear function is to the power of 1.

OpenStudy (anonymous):

oka.. so its B right?

OpenStudy (anonymous):

x^2 is to the power of 2 x^1 is to the power of 1 but we don't actually write the power to 1 but rather just x.

OpenStudy (anonymous):

nope, not b because it has x^2 so it's to the power of 2.

OpenStudy (anonymous):

OH OKA.. HOLD ON LET ME SEE WHICH ONE IT IS SO ITS IN CAPS FORGOT TO TAKE OFF CAPS

OpenStudy (anonymous):

so the only one left with out a power of 2 is d.) nno i think im still wrong

OpenStudy (anonymous):

hello anybody here?

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is d.

OpenStudy (anonymous):

Ok so my last option would be d

OpenStudy (anonymous):

thats what me thought lol skittles luves you lol

OpenStudy (anonymous):

x+2=y you can see the power of x is 1 because x^1 is the same thing as x.

OpenStudy (anonymous):

even if you rearrange it into y=-x+2 the form of y=ax+b you can see the power of x is still one, therefore it is linear.

OpenStudy (anonymous):

and like JayDs said you would be looking for a constant slope with a degree of x.

OpenStudy (anonymous):

thank yall an js my name is skittles an i luv you guys for helpin me